### v1

# On the Application of Model Predictive Control to a Weighted Coverage Path Planning Problem

Abstract
This paper considers the application of Model Predictive Control (MPC) to a weighted coverage path planning (WCPP) problem. The problem appears in a wide range of practical applications, such as search and rescue (SAR) missions. The basic setup is that one (or multiple) agents can move around a given search space and collect rewards from a given spatial distribution. Unlike an artificial potential field, each reward can only be collected once. In contrast to a Traveling Salesman Problem (TSP), the agent moves in a continuous space. Moreover, he is not obliged to cover all locations and/or may return to previously visited locations. The WCPP problem is tackled by a new Model Predictive Control (MPC) formulation with so-called Coverage Constraints (CCs). It is shown that the solution becomes more effective if the solver is initialized with a TSP-based heuristic. With and without this initialization, the proposed MPC approach clearly outperforms a naive MPC formulation, as demonstrated in a small simulation study.
I Introduction
Many path planning applications in robotics try to find an optimal path while maximizing some sort of reward. The reward is commonly related to the proximity to given reference value, guiding the system via an artificial potential field (AFP), or keeping it away from unsafe areas via barrier functions. Model Predictive Control (MPC) casts this problem into a numerical optimization program, with an objective function and constraints. Over the past years, it has become a standard approach for path planning, due to its natural handling of general reward functions, system dynamics, and input and state constraints.
Coverage Path Planning (CPP) aims for the system to cover an entire area of the state space, or as large a part of it as possible. Problem instances appear in many robotic applications, such autonomous lawn mowers, vacuum cleaners, agricultural robots, and arial / underwater drones used for inspection or surveillance. From the perspective of common MPC-based path planning, this means a uniform objective function. However, for CPP, the objective function changes dynamically in the sense that the reward at an already visited position, and some radius around it, drops to zero.
Efficient motion patterns or policies are commonly used for CPP, including a boustrophedon (snake-like) path or straight driving with random reflection angles when hitting boundaries [1]. The main principle is, clearly, to minimize any overlaps in the track of the robot.
This paper considers the extended problem of Weighted Coverage Path Planning (WCPP). The main difference to the CPP is that the objective function is not uniform, but weighted by a coverage priority. Problem instances of this also appear frequently, e.g., in search and rescue (SAR) or surveillance missions, where the goal is to find or detect a target whose probability of presence on a map is not uniformly distributed. Correspondingly, the motivational example for this paper is an unmanned aerial vehicle (UAV) with the task of finding a missing person in a given area as fast as possible. The coverage priority is indicated by a probability map, which serves as the reward function and represents a belief distribution integrating all available information at the current time, e.g., from sensor measurements, the observation of eye witnesses, or human behavioral models in the given map [2]. For practical applications, a probabilistic model could be used to dynamically update the belief distribution based on all available information, and thus guide the planning of the mission [3].
I-A Existing Literature
CPP is about finding a path that covers all specified points or an entire area or region of interest while avoiding obstacles [4]. A comprehensive overview of CPP algorithms for robotic applications is given in [5]. The article covers both classical and heuristic-based algorithms. These two categories include a whole variety of basic approaches, such as AFPs, greedy search and graph search algorithms, and bio-inspired approaches, such as genetic algorithms or ant colony optimization.
A key point in the discussion in [5] is that the resulting paths should be as smooth as possible. Namely, the avoidance of sharp turns in the path prevent premature wear of the robot’s components and it increases the efficiency, especially in UAV applications, where it advantageous for the average speed of the UAV to fly straight ahead or in smooth curves. In this spirit, the approach in [6] proposes a path search algorithm using ant colony optimization based on the Lin-Kernighan heuristic, followed by a smoothing step using a customized approach based on Fourier series. The resulting dynamically smooth trajectory is then fed to the UAV, which is operated by an MPC-based controller.
Numerical optimization has become increasingly popular for path planning over the recent years, due to the availability of more powerful hardware and increasingly efficient solvers. MPC, in particular, has been successfully employed for collision-free path planning for autonomous road vehicles [7] and for UAVs [8]. MPC has been used in combination with AFPs [9, 10], and also for CPP with obstacles using a mixed-integer linear programming (MILP) formulation [11]. The area is covered by a uniform grid, where each cell defines a single way point. The way points are subsequently represented as discrete decision variables within the optimization problem, where the objective is to cover the maximum number of (equally weighted) way points.
Related problems to the CPP include the Traveling Salesman Problem (TSP) and its variants, most notably the Orienteering Problem (OP) [12]. In contrast to the TSP, where all vertices must be visited and the goal is to minimize the traveled distance, the OP is concerned with maximizing the total reward collected within a limited time frame, without necessarily visiting all the vertices [13]. Both of these problems are known to be NP-hard [14]. Yet it is desirable, for many applications, to extend them further by including the dynamics of an agent in the problem formulation [15]. Recently, a mathematical framework to tackle this problem using techniques from optimal control has been proposed [15, 16]. In addition to the dynamics of an agent, this formulation also accounts for the movement of the sensors, i.e., in this case, a camera. However, the approach relies on nonsmooth calculus and considers only discrete regions of interest.
I-B Contributions
This paper proposes a new MPC approach for the WCPP problem, using coverage constraints (CCs). The proposed MPC formulation works with any agent moving governed by a nonlinear dynamic model and moving in a continuous space with obstacles.
The reward function is arbitrary, but in contrast to an AFP, each reward can only be collected once. This is enforced by the use of quadratic constraints. In contrast to the TSP, the agent is not obliged (and may in fact be far from able to) cover all locations within the given prediction horizon.
Furthermore, the paper shows that the solution becomes more effective if the MPC solver is initialized with a TSP-based heuristic. The heuristic is based on a set of key points, which are derived from a Gaussian mixture model that is used to approximate the reward function.
II Methods
Consider a reward function $r:\mathbb{R}^{n}\to\mathbb{R}$, which maps a state $x\in\mathbb{R}^{n}$ to its reward $r(x)$. It is assumed that $r$ is twice continously differentiable, in order to enable gradient-based optimization. The goal is to find a trajectory $\bar{h}:[t_{0},t_{1}]\to\mathbb{R}^{n}$ with $\bar{h}(t_{0})=x_{0}\in\mathbb{R}^{n}$ satisfying the dynamic constraints $\dot{x}=\bar{f}(x,u)$ of an agent. The objective is to maximize the reward along the path by choosing the control input $u\in\mathbb{R}^{m}$ between time $t_{0}$ and $t_{1}$. Specifically, we seek to maximize the integral
$\int_{t_{0}}^{t_{1}}r\left(\bar{h}(t)\right)\mathrm{d}t$ (1)
As shown in Section II-A, the objective (1) can be easily incorporated into the objective function of the MPC. However, as the problem is highly non-linear in the input signal $u(t)$, we try to improve the solution by providing a good initial guess to the NLP solver. The problem is therefore solved in a hierachical fashion, consisting of multiple steps:
  1. 1.
    Find a set of key points using Gaussian mixture models as described in Section II-B;
  2. 2.
    Find an optimal path going through all key points using a travelling salesman formulation as described in Section II-B;
  3. 3.
    Solve the receding horizon optimal control problem as explained in Section II-A, using the solution obtained in the previous step as an initial guess.
Note that in the case of a discrete probability grid, the continuous reward function can be constructed based on a B-spline interpolation, for example.
II-A Finding a trajectory with MPC
We consider a discrete-time optimal control problem of finding the optimal sequence of control inputs $u_{0},u_{1},…,u_{N-1}\in\mathbb{R}^{m}$, driving the states $x_{0},x_{1},…,x_{N}\in\mathbb{R}^{n}$ of a dynamical system $x_{k+1}=f(x_{k},u_{k})$ for $k=0,1,…,N-1$, where $N$ is the horizon length.
[…]
$J(\mathbf{x}_{0:N},\mathbf{u}_{0:N-1})=c_{1}\sum_{k=0}^{N-1}u_{k}^{\mathsf{T}}Ru_{k}-c_{2}\sum_{k=0}^{N}r(x_{k})\;.$ (2)
For the control input costs we choose a quadratic formulation, defined by a positive definite cost matrix $R\in\mathbb{R}^{m\times m}$. The second term in the objective function maximizes the reward along the path, approximating the integral in (1). Both terms are weighted by the (positive) parameters $c_{1}$ and $c_{2}$, respectively.
The constraints (II-A) and (II-A) keep the states and inputs in the set of admissible states (the region of the map) $\mathbb{X}$ and in the set of admissible inputs (accelerations) $\mathbb{U}$, respectively. The constraint (II-A) sets the initial state of the agent.
To prevent the agent from just collecting the same reward multiple times, we introduce additional coverage constraints
$\rho(x_{i},x_{j})\geq V-\epsilon_{i},\quad i=1,…,N,~{}j=0,…,i-1$ (3)
for some distance metric $\rho(x_{i},x_{j})$. These constraints force the agent to move by ensuring that each state $x_{i}$ along the trajectory is at least distance $V$ away from each previous state $x_{j}$ for $j<i$. The parameter $V$ is related to the visibility of the agent, i.e., it determines how far the agent needs to travel in order to collect a new reward. Note that the constraint is implemented as a soft constraint with the slack variables $\epsilon_{i}\in\mathbb{R}$ in order to allow the agent to slightly violate a CC if this leads to a better trajectory. The slack variables appear as an additional term in the modified objective function
$J(\mathbf{x}_{0:N},\mathbf{u}_{0:N-1})=\\ c_{1}\sum_{k=0}^{N-1}u_{k}^{\mathsf{T}}Ru_{k}-c_{2}\sum_{k=0}^{N}r(x_{k})+c_{3}\sum_{k=0}^{N}\epsilon_{k}\;.$ (4)
Note that the dynamics (II-A) in the MPC formulation are generally arbitrary. For the implementation, simple linear dynamics $f(x_{k},u_{k})=Ax_{k}+Bu_{k}$ are assumed, in the form of a double integrator:
$A=\begin{bmatrix}1&0&\Delta t&0\\ 0&1&0&\Delta t\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},~{}B=\begin{bmatrix}0&0\\ 0&0\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}$
. Here, the states are the positions and velocities, and the inputs are the accelerations. The function $\rho(x_{i},y_{i})$ is defined as the distance between the positions of two states, i.e.,
$\rho(x_{i},x_{j})=\sqrt{(x_{i,1}-x_{j,1})^{2}+(x_{i,2}-x_{j,2})^{2}}\;.$ (5)
This leads to the quadratic CCs
$(x_{i,1}-x_{j,1})^{2}+(x_{i,2}-x_{j,2})^{2}\geq V^{2}-\epsilon_{i}\\ \forall\;i=1,…,N\;\text{and}\;j=0,…,i-1\;.$ (6)
Figure 1: Solution of the MPC for the same scenario as in Figure 5(a), but without an initial guess.
II-B Finding an initial solution with a TSP based heuristic
Depending on the particular problem instance, the MPC with CCs may itself already yield satisfactory results for the WCPP problem. However, it can be observed that in some cases the solution gets stuck in a local minimum, leading to a sub-optimal behavior. An example of this is shown in Figure 1. It can be observed that the effect strongly depends on the initial solution given to the NLP solver. Therefore, heuristic is employed to remedy this problem, as explained in this section.
The idea of the heuristic is to find a set of key points, each of which is likely to yield a high reward. Additionally, key points should approximate the global structure of the reward function. A good initial solution can then be obtained by finding a path that passes through all the key points. This allows the global structure to be exploited, rather than just optimizing locally as would be the case if the MPC were simply run without this step.
The presented algorithm is based on Gaussian mixture models (GMMs). They can approximate any probability density function [17] and have also been used before to prioritize search subregions in a probability map [18]. Note that it is not necessary to consider the reward in (1) as a probability density function. However, we make the additional assumptions that the reward is always positive and that the total collectable reward is finite:
$r(x)\geq 0\quad\forall x\in\mathbb{R}^{n}\;,$ (7a)
$\int r(x)\mathrm{d}x<\infty\;.$ (7b)
Note that under assumptions (7a) and (7b), any reward function can be considered a probability density function (up to a scaling factor).
A Gaussian mixture with $n$ components is a function given by
$p(x)=\sum_{k=1}^{n}w_{k}\mathcal{N}(x;\mu_{k},\Sigma_{k})\;,$ (8)
where $\mathcal{N}(x;\mu_{k},\Sigma_{k})$ is a multivariate normal distribution and $w_{k}$ are the weights of the components,] and $n$ denotes the number of mixture components, i.e., the number of key points.
The parameters $w_{k},\mu_{k},\Sigma_{k}$ are optimized using the Expectation Maximization (EM) algorithm [19]. Hence, by limiting the number of mixture components, the centers of the Gaussians are expected to be at positions yielding a high reward. It can lead to better results if only the best (in terms of mean value) $m<n$ components are used, for instance $m=n/2$. In this case, we choose $m$ to be the number of key points. See Figure 2 for results using this approach.
Figure 2: Key points found using Gaussian mixtures with $n=20$ ($m=10$) and $n=40$ ($m=20$) components respectively.
Given the set of key points, we now want to find the shortest path that passes all of them. However, this problem is hard in itself, since even finding a shortest path through all key points without considering rewards is an instance of the TSP. Although there are variants of TSP that also take into account the potential rewards for each city, the regular variant is used in the following.
Typically, the TSP aims for a route returning back to the origin, which is not the desired outcome in this case. To circumvent this issue, one can set the cost of travelling from any location to the current location to zero. Assuming $p_{0}$ is the current location, the cost function for two points $p_{i}$ and $p_{j}$ is given as
$d(p_{i},p_{j})=\begin{cases}0&\text{ if }j=0\\ \|p_{i}-p_{j}\|_{2}&\text{ otherwise}\end{cases}\;.$
This forces the algorithm to start at the current position and returns the desired result. The implementation uses a Dantzig–Fulkerson–Johnson formulation and is based on [20, 21]. See Figure 3 for results.
To feed the travelling salesman tour from the as an initial solution to the MPC problem (II-A), the tour is discretized using the maximum velocity as the step size. Figure 4 shows the result of running the MPC as the closed-loop controller, given the tour from Figure 3 as the initial solution.
Figure 3: A travelling salesman tour through all key points (orange), starting from the initial position $x_{0}$ (red). The tour also ends in $x_{0}$, which is ommitted here.
Figure 4: The resulting open-loop trajectory of the MPC controller based on the initial solution obtained from the tour in Figure 3.
III Results
The methods described in Section II have been implemented in Julia, using the JuMP toolkit [22] for defining optimization problems. As the underlying solver, HiGHS [23] has been used for solving mixed-integer problems (TSP), and Ipopt [24] for non-linear continuous optimization problems (MPC). The parameters of the objective function (II-A) have been selected as $c_{1}=1,c_{2}=1000,c_{3}=100$. All experiments have been performed on a machine with an Intel i5-1350P processor.
The implementation has been tested for scenarios in which a path is calculated for a UAV moving in a probability map, with the objective being to move through areas with the highest probability. Three different scenarios with different horizon lengths have been considered. The first two scenarios in Figure 5(a) and Figure 5(b) are based on a given discrete probability map with sizes $600\times 600$ meters and $800\times 800$ meters, respectively. Both maps are defined on a regular grid with a cell size of $25$ meters. The initial position can be chosen freely. To obtain a continuous reward function from the discrete probability map, a B-spline interpolation is used. The third scenario in Figure 5(d) is defined by combining multiple Gaussian distributions.
The runtime performance for all three scenarios is shown in Table I. As can be seen in the table, most of the time is spent on solving the nonlinear optimization problem. Nevertheless, the quality of the solution compared to running the MPC without an initial guess could be improved in all secenarios. This can be observed in Figure 1, which shows the solution for Scenario 1 but without an initial guess. The total reward value is lower, and the controller tends to stay in a local region instead of going to regions with high rewards.
(a) Scenario 1: $600\times 600$ map
with horizon length $N=50$ (left) and $N=100$ (right)
(b) Scenario 2a: $800\times 800$ map
with horizon length $N=50$ (left) and $N=100$ (right)
(c) Scenario 2b: $800\times 800$ map
with horizon length $N=100$ (left) and $N=150$ (right)
(d) Scenario 3: $800\times 800$ map with horizon length $N=75$ (left) and $N=150$ (right)
Figure 5: Results for three different scenarios, with increased horizon length on the right side. The red marker denotes the initial position of the UAV.
$n$ $N$ GMM TSP MPC Total
Scenario 1 20 50 0.18s 0.004s 6.42s 6.6s
Scenario 1 20 100 0.23s 0.006s 28.39s 28.63s
Scenario 2a 40 50 0.41s 0.072s 3.31s 3.79s
Scenario 2a 40 100 0.51s 0.27s 38.15s 38.93s
Scenario 2b 40 100 0.4s 0.03s 16.43s 16.86s
Scenario 2b 40 150 0.4s 0.049s 53.9s 54.35s
Scenario 3 40 75 0.37s 0.007s 4.1s 4.48s
Scenario 3 40 150 0.62s 0.012s 41.89s 41.52s
TABLE I: Runtime performance of the scenarios in Figure 5, where $n$ is the number of mixture components, and $N$ is the horizon length of the MPC
IV Conclusion
The paper presents a new approach of using MPC for the WCPP problem, where a spatially distributed reward is collected by an agent with continuous dynamics. To this end, the new concept of CCs is introduced.
The performance of the solution is improved by providing an initial guess based on the solution of a TSP, which finds the optimal path going through a set of key points obtained by approximating the reward function using GMMs.
Specifically, we consider path planning for SAR missions, where the reward is given as the probability of finding a missing person in a given area or map. Correspondingly, the proposed approach has been evaluated on a number of specific problem instances, demonstrating its practical and computational viability.
Ross et al. showed that TSP problems can also be solved using optimal control formulations [16], so it may be possible to merge the steps from Sections II-B into a single formulation in a future work. Furthermore, it could be interesting to try other MPC formulations and/or CCs and evaluate them in the context of the WCPP.